One type of candy sells for $2 per
pound, and another type sells for $3 per pound. An
order for 18 pounds of candy costs $47. How much of
each type of candy was bought?

Answer by edjones(8007)

Let x=lbs that sell for $2 and y=lbs that sell for $3.
x + y = 18
2x+3y=47
.
2x+2y=36 Multiply each side by 2.
2x+3y=47
--------subtract
-y=-11
.
y=11 lbs
x=7 lbs
.
Ed